show steps on how to break down and graph the equation as a line 2x-5y=-20
Answer by AAfter Search(61)

2x-5y=-20
=> 2x + 20 = 5y
=> y = (2x + 20)/5 = 2x/5 + 4
For x = 0, y = 4,
For x = 5, y = (2 x 5)/5 + 4 = 2 + 4 = 6
Hence, the two points to draw the line are (0,4) and (5,6).